# Oho 1-1, Tsukuba, Ibaraki 305-0801, Japan

### v1

@@ Abstract
Recent advancements in deep learning models have significantly enhanced jet classification performance by analyzing low-level features (LLFs). However, this approach often leads to less interpretable models, emphasizing the need to understand the decision-making process and to identify the high-level features (HLFs) crucial for explaining jet classification. To address this, we introduce an analysis model (AM) that analyzes selected HLFs designed to capture important features of top jets. Our AM mainly consists of the following three modules: a relation network analyzing two-point energy correlations, mathematical morphology and Minkowski functionals for generalizing jet constituent multiplicities, and a recursive neural network analyzing subjet constituent multiplicity to enhance sensitivity to subjet color charges. We demonstrate that our AM achieves performance comparable to the Particle Transformer (ParT) in top jet tagging and simulated jet comparison at the hadronic calorimeter angular resolution scale while requiring fewer computational resources. Furthermore, as a more constrained architecture than ParT, the AM exhibits smaller training uncertainties because of the bias-variance tradeoff. We also compare the information content of AM and ParT by decorrelating the features already learned by AM.
@@ 1 Introduction
What if HLF-based taggers are competitive with more general taggers such as graph neural networks? There would be two immediate advantages: network predictions would have less training uncertainty, and simpler taggers would require fewer computational resources. In principle, if we could design a set of HLFs that fully capture features crucial for jet tagging, we could develop a high-performing jet tagger based on HLFs. Additionally, such taggers could be further utilized for calibrating simulations when they are trained to distinguish simulations from the experimental data with high precision.
@@
In Section 4, we estimate the statistical significance of the difference between our analysis model and ParT using the bootstrap method. Our model exhibits smaller training uncertainties while maintaining performance comparable to ParT, within the range of $2\sim 3\sigma$. We also assess the difference in information learned by the two models by penalizing information in ParT that has already been learned by the analysis model. We summarize our findings and conclude in Section 5.
@@ 2.1 Anatomy of Top Jets and Feature-wise Analysis Modules
and it will be trained by minimizing the binary cross-entropy loss function with the prior of each class to 0.5.
@@ 2.2 Global Features: Basic Jet Kinematics
We first select the global features $x_{\mathrm{kin}}$ mainly to provide characteristic angular scales of jets, which are given by the ratio of jet mass $m$ and transverse momentum $p_{T}$ to all the analysis modules. In particular, we use the kinematical variables of the whole jet $\mathrm{\mathbf{J}}$ and the trimmed jet $\mathrm{\mathbf{J}}_{\mathrm{trim}}$to capture the angular scale of top quark decay. [^2] For jet trimming, we use $k_{T}$ algorithm Catani:1993hr ; Ellis:1993tq with radius 0.2 and keep subjets whose energy fraction is larger than 0.05. We additionally provide the variables of the leading $p_{T}$ subjet $\mathrm{\mathbf{J}}_{\mathrm{lead}}$ to implicitly capture the W-boson decay scale. [^3] The leading $p_{T}$ subjet is the highest $p_{T}$ anti-$k_{T}$ subjet Cacciari:2008gp with radius 0.2. The resulting global features consist of the following inputs:
@@ 2.4 Constituent Multiplicity Distributions and Mathematical Morphology
Figure 4: Average counts of the jet constituents per $p_{T}$ bins for top and QCD jet samples. The jet constituents are Delphes Eflow objects of QCD or top jets pixellated with the size $\Delta=0.1\times 0.1$. The bin boundaries are at $p_{T}=[0.1$, $0.5\times(\sqrt{2})^{k}$ ($k=0$,..), $512$] GeV. See Section 3.1 for our event sample and pre-processing details.
@@
The differences in expectations between the QCD and top jets are more significant than the above systematic uncertainties between the simulations. In the low $p_{T}$ region, QCD jets typically have more constituents than top jets due to gluon (sub)jets and the high mass selection. On the other hand, the high $p_{T}$ region captures features of the hard process. The top jet histogram notably shows a prominent bump at $\sim 50$ GeV. This bump is mostly due to the subjets originating from the $W$ boson and the hadronization of the corresponding boosted $q\bar{q}$ system. This multiplicity distribution conditioned on $p_{T}$ captures the differences between top and QCD jets and the systematic differences between simulations. Including this histogram and its generalization for the AM will significantly enhance the information encoding performance.
@@
These MFs are used with dilation operation to encode non-trivial changes in the topology and geometry appearing at various scales, similar to persistence homology in topological data analysis 10.3389/frai.2021.667963 . Let us define a dilated image $\mathcal{I}(R)$ as a Minkowski sum of the set of constituents $\mathcal{C}$ and a disk $B(R)$ with radius $R$, i.e.,444 Note that in our previous works Chakraborty:2020yfc ; Lim:2020igi , we constructed a model using the MFs dilation by a square and compared the results with CNN—both of the models process jet images with Manhattan geometry. In this paper, we use the disk respecting the rotational symmetry of Euclidean geometry instead. The edge variable of ParT is invariant under rotation, and this change aligns persistence analysis through the MF module consistent with ParT.
@@
However, as soon as the disks come in contact, this relationship no longer holds, indicating that a non-trivial topological change occurred at the dilation scale. The MFs after the dilations are computed using shapely package Gillies_Shapely_2023 based on GEOS library libgeos . For the numerical evaluations, the disks are approximated by regular 64-sided polygons.
@@
Note that the diameter $2R$ is the parameter describing the correlation scale in this persistence analysis. The first MFs with $R=0.025$ are simply the constituent multiplicities since we are using angular resolution $0.1$ and no disks are overlapping each other; the MFs will be simply given as Eq. 11. For $k>1$, some disks may overlap, and MFs can capture non-trivial topological changes during the dilation. For $R>0.2$, we switch to a coarser grid. Most small-scale effects are smeared out at this scale, leaving only large-scale information less sensitive to the $R$ choice, such as hard substructure and large-angle radiations. We limit the angular scale up to $R=0.3$ in this paper.
@@ 2.5 Recursive Neural Network for Subjet Constituent Multiplicities and Kinematics
Here, $p_{k}$ is the subjet 4-momentum, and $N_{c,k}$ denotes the number of constituents in the subjet. For simplicity, we do not use the full Minkowski functionals; the $N_{c,k}$ alone provides sufficient information when this module operates together with the others.
@@ 3.2 Particle Transformer
Since our aim is to compare classifier performance with inputs having a finite angular resolution $0.1$, we will mask out small-scale information by utilizing jet images Cogan:2014oua ; Almeida:2015jua ; deOliveira:2015xxd . The jet constituents are first preprocessed as described in Lim:2020igi . This preprocessing involves shifting, rotating, and flipping on the $(\eta,\phi)$ coordinate to ensure that the leading subjet is at the origin, the second subjet lies on the positive $\phi$-axis ($\eta=0$ and $\phi>0$), and the third subjet is on the right side of the plane ($\eta>0$). After the preprocessing, the jet constituents are pixelated into bins of size $\Delta\eta=\Delta\phi=0.1$. Pixels with non-zero energy deposits are converted into massless 4-momenta using their $(\eta,\phi)$ coordinates and energy deposits. This set of 4-momenta is used as ParT inputs.
@@ 3.3.1 Comparing Classification Performance
A) $x_{\rm kin}$, $x_{S_{2}}$ 0.936 0.919 0.571 0.569
B) $x_{\rm kin}$, $x_{\mathrm{subj}}$ 0.938 0.922 0.575 0.578
C) $x_{\rm kin}$, $x_{S_{2}}$, $x_{\mathrm{subj}}$ 0.941 0.925 0.579 0.585
@@ 4 Statistical Significance of the Difference between Analysis Model and Particle Transformer
ParT-best (as I) 0.5920$\pm$0.004 2.700$\pm$0.040(±$\pm 0.006$) 5.20$\pm$0.10(±$\pm 0.02$)
ParT-2nd (as II) 0.5910$\pm$0.004 2.700$\pm$0.040(±$\pm 0.006$) 5.14$\pm$0.08(±$\pm 0.02$)
@@ 4.1 Estimating Uncertainty of Performance Metrics by Bootstrapping
We estimate uncertainties associated with network training by using the bootstrap method. This method involves the generation of bootstrap datasets from the original dataset by sampling with replacement. In this resampling procedure, we treat the empirical distribution of the training dataset as an approximate true distribution. The bootstrap datasets effectively act as representative samples of this approximate true distribution. By training networks again on each of these datasets, this bootstrap method allows us to estimate the variance of training results. For this study, we use ten bootstrap datasets. Note that our AM is computationally cheaper than ParT, so this bootstrap analysis can be done much faster. The computational resources utilized by AM and ParT are explained in Appendix B.
@@
Figure 6: Distribution of the difference between the classifier outputs of the best model and the second best model for $\mathrm{P_{Q}}$ vs. $\mathrm{H_{Q}}$ classification. The red solid line is for ParT, and the red dot-dashed line is for AM. The distribution for ParT is significantly broader than that of AM. Additionally, the blue dotted line is for the AM-PIP with only the morphology analysis module active, and the blue dashed line is for the AM-PIP with relation network and subjet analysis modules active. These two blue lines are for demonstrating output variations in simpler models compared to the full AM.
@@ 4.2 Comparing the Information Content of Classifiers using Odds Ratio
An odds ratio greater than 1 indicates that the classifier $\mathrm{I}$ is more confident in its signal prediction by the factor of the odds ratio than the classifier $\mathrm{II}$. If both models are identical, the odds ratio will be 1.777Note that this reference point 1 may change if you compare classifiers trained using different class priors. Here, we only consider classifiers trained using class prior probability to 0.5. Therefore, deviations from 1 in the odds ratio indicate differences in the predictions of the two classifiers for a given sample.
@@
A more detailed derivation of this interpretation is described in Appendix C.

### v2

@@ Abstract
Recent advancements in deep learning models have significantly enhanced jet classification performance by analyzing low-level features (LLFs). However, this approach often leads to less interpretable models, emphasizing the need to understand the decision-making process and to identify the high-level features (HLFs) crucial for explaining jet classification. To address this, we consider the top jet tagging problems and introduce an analysis model (AM) that analyzes selected HLFs designed to capture important features of top jets. Our AM mainly consists of the following three modules: a relation network analyzing two-point energy correlations, mathematical morphology and Minkowski functionals for generalizing jet constituent multiplicities, and a recursive neural network analyzing subjet constituent multiplicity to enhance sensitivity to subjet color charges. We demonstrate that our AM achieves performance comparable to the Particle Transformer (ParT) while requiring fewer computational resources in a comparison of top jet tagging using jets simulated at the hadronic calorimeter angular resolution scale. Furthermore, as a more constrained architecture than ParT, the AM exhibits smaller training uncertainties because of the bias-variance tradeoff. We also compare the information content of AM and ParT by decorrelating the features already learned by AM. Lastly, we briefly comment on the results of AM with finer angular resolution inputs.
@@ 1 Introduction
What if HLF-based taggers are competitive with more general taggers such as graph neural networks? There would be two immediate advantages: network predictions would have less training uncertainty, and simpler taggers would require fewer computational resources. In principle, if we could design a set of HLFs that fully capture features crucial for jet tagging, we could develop a high-performing jet tagger based on HLFs. Inputs used by such taggers highlight the distribution that should be calibrated by the data or give us the clear benchmarks that event simulation should satisfy.
@@
In Section 4, we estimate the statistical significance of the difference between our analysis model and ParT using the bootstrap method. Our model exhibits smaller training uncertainties while maintaining performance comparable to ParT, within the range of $2\sim 3\sigma$. We also assess the difference in information learned by the two models by penalizing information in ParT that has already been learned by the analysis model.
Section 5 discusses the potential high-level features that start appearing below the hadronic calorimeter scale and presents the classification performance of AM tweaked for the electromagnetic calorimeter scale analysis without accounting for the additional high-level features. We summarize our findings and conclude in Section 6.
@@ 2.1 Anatomy of Top Jets and Feature-wise Analysis Modules
and it will be trained by minimizing the binary cross-entropy loss function.  [^2] For training, we use the same number of data sets to make the prior of each class to be 0.5.
@@ 2.2 Global Features: Basic Jet Kinematics
We first select the global features $x_{\mathrm{kin}}$ mainly to provide all the analysis modules with the characteristic angular scales of jets, which are given in terms of the jet mass $m$ and transverse momentum $p_{T}$. In particular, we use the kinematical variables of the whole jet $\mathrm{\mathbf{J}}$ and the trimmed jet $\mathrm{\mathbf{J}}_{\mathrm{trim}}$ to capture the angular scale of top quark decay. [^3] For jet trimming, we use $k_{T}$ algorithm Catani:1993hr ; Ellis:1993tq with radius 0.2 and keep subjets whose energy fraction is larger than 0.05. We additionally provide the variables of the leading $p_{T}$ subjet $\mathrm{\mathbf{J}}_{\mathrm{lead}}$ to implicitly capture the W-boson decay scale. [^4] The leading $p_{T}$ subjet is the highest $p_{T}$ anti-$k_{T}$ subjet Cacciari:2008gp with radius 0.2. The resulting global features consist of the following inputs:
@@ 2.4 Constituent Multiplicity Distributions and Mathematical Morphology
[…]
@@
The differences in expectations between the QCD and top jets are more significant than the above systematic uncertainties between the simulations. In the low $p_{T}$ region, QCD jets typically have more constituents than top jets due to gluon (sub)jets and the high mass selection. On the other hand, the high $p_{T}$ region captures features of the hard process. The top jet histogram notably shows a prominent bump at $\sim 50$ GeV. This bump is mostly due to the subjets originating from the $W$ boson, and the hadronization of the corresponding boosted $q\bar{q}$ system. This multiplicity distribution conditioned on $p_{T}$ captures the differences between top and QCD jets and the systematic differences between simulations. Including this histogram and its generalization for the AM will significantly enhance the information encoding performance.
@@
These MFs are used with dilation operation to encode non-trivial changes in the topology and geometry appearing at various scales, similar to persistence homology in topological data analysis 10.3389/frai.2021.667963 . Let us define a dilated image $\mathcal{I}(R)$ as a Minkowski sum of the set of constituents $\mathcal{C}$ and a disk $B(R)$ with radius $R$, i.e.,555 Note that in our previous works Chakraborty:2020yfc ; Lim:2020igi , we constructed a model using the MFs dilation by a square and compared the results with CNN—both of the models process jet images with Manhattan geometry. In this paper, we use the disk respecting the rotational symmetry of Euclidean geometry instead. The edge variable of ParT is invariant under rotation, and this change aligns persistence analysis through the MF module consistent with ParT.
@@
However, as soon as the disks come in contact, this relationship no longer holds, indicating that a non-trivial topological change occurred at the dilation scale. The MFs after the dilations are computed using shapely package Gillies_Shapely_2023 based on GEOS library libgeos . For the numerical evaluations, the disks are approximated by regular 64-sided polygons. The Python code used for computing the MFs is available at https://github.com/sunghak-lim/mfjet.
@@
Note that the diameter $2R$ is the parameter describing the correlation scale in this persistence analysis. The first MFs with $R=0.025$ are simply the constituent multiplicities since we are using angular resolution $0.1$ and no disks are overlapping each other; the MFs will be simply given as Eq. 11. For $k>1$, some disks may overlap, and MFs can capture non-trivial topological changes during the dilation. For $R>0.2$, we switch to a coarser grid. Most small-scale effects are smeared out at this scale, leaving only large-scale information less sensitive to the $R$ choice, such as hard substructure and large-angle radiations. We limit the angular scale up to $R_{\mathrm{max}}=0.3$ in this paper. The dependence on the choice of $R_{\mathrm{max}}$ can be found in Section B.
@@ 2.5 Recursive Neural Network for Subjet Constituent Multiplicities and Kinematics
Here, $p_{k}$ is the subjet $p_{T}$, $\eta$, $\phi$ and $m$, and $N_{c,k}$ denotes the number of constituents in the subjet. For simplicity, we do not use the full Minkowski functionals; the $N_{c,k}$ alone provides sufficient information when this module operates together with the others.
@@ 3.2 Particle Transformer
Since our aim is to compare classifier performance with inputs having a finite angular resolution $0.1$, we will mask out small-scale information by utilizing jet images Cogan:2014oua ; Almeida:2015jua ; deOliveira:2015xxd . The jet constituents are first preprocessed as described in Lim:2020igi . This preprocessing involves shifting, rotating, and flipping on the $(\eta,\phi)$ coordinate to ensure that the leading subjet is at the origin, the second subjet lies on the positive $\phi$-axis ($\eta=0$ and $\phi>0$), and the third subjet is on the right side of the plane ($\eta>0$). After the preprocessing, the jet constituents are pixelated into bins of size $\Delta\eta=\Delta\phi=0.1$. Pixels with non-zero energy deposits are converted into massless 4-momenta using their $(\eta,\phi)$ coordinates and energy deposits. This set of 4-momenta is used as ParT inputs. The results using a finer resolution of $\Delta\eta=\Delta\phi=0.025$ will be shown in Section 5.
@@ 3.3.1 Comparing Classification Performance
A) $x_{\rm kin}$, $x_{S_{2}}$ 0.936 0.919 0.571 0.5690
B) $x_{\rm kin}$, $x_{\mathrm{subj}}$ 0.938 0.922 0.575 0.5780
C) $x_{\rm kin}$, $x_{S_{2}}$, $x_{\mathrm{subj}}$ 0.941 0.925 0.579 0.5850
@@ 4 Statistical Significance of the Difference between Analysis Model and Particle Transformer
ParT-best (as I) 0.5920$\pm$0.0040 2.700$\pm$0.040(±$\pm 0.006$) 5.20$\pm$0.10(±$\pm 0.02$)
ParT-2nd (as II) 0.5910$\pm$0.0040 2.700$\pm$0.040(±$\pm 0.006$) 5.14$\pm$0.08(±$\pm 0.02$)
@@ 4.1 Estimating Uncertainty of Performance Metrics by Bootstrapping
We estimate uncertainties associated with network training by using the bootstrap method. This method involves the generation of bootstrap datasets from the original dataset by sampling with replacement. In this resampling procedure, we treat the empirical distribution of the training dataset as an approximate true distribution. The bootstrap datasets effectively act as representative samples of this approximate true distribution. By training networks again on each of these datasets, this bootstrap method allows us to estimate the variance of training results. For this study, we use ten bootstrap datasets. Note that our AM is computationally cheaper than ParT, so this bootstrap analysis can be done much faster. The computational resources utilized by AM and ParT are explained in Appendix C.
@@
Figure 6: Distribution of the difference between the classifier outputs of the best model and the second best model for $\mathrm{P_{Q}}$ vs. $\mathrm{H_{Q}}$ classification. The red solid line is for ParT, and the blue dot-dashed line is for AM. The distribution for ParT is significantly broader than that of AM. Additionally, the green dotted line is for the AM-PIP with only the morphology analysis module active, and the orange dashed line is for the AM-PIP with relation network and subjet analysis modules active. These two extra lines are for demonstrating output variations in simpler models compared to the full AM.
@@ 4.2 Comparing the Information Content of Classifiers using Odds Ratio
An odds ratio greater than 1 indicates that the classifier $\mathrm{I}$ is more confident in its signal prediction by the factor of the odds ratio than the classifier $\mathrm{II}$. If both models are identical, the odds ratio will be 1.888Note that this reference point 1 may change if you compare classifiers trained using different class priors. Here, we only consider classifiers trained using class prior probability 0.5. Therefore, deviations from 1 in the odds ratio indicate differences in the predictions of the two classifiers for a given sample.
@@
A more detailed derivation of this interpretation is described in Appendix D.
@@
5 Discussion on Analysis Model and High-Level Features at Sub-Hadronic Calorimeter Angular Resolution
Throughout the previous analyses, we used datasets pixelated or binned to the angular resolution of the hadronic calorimeter (HCAL), $\Delta R_{H}\simeq 0.1$. The classification performance of our AM is consistent with one of the state-of-the-art models, Particle Transformer. A natural follow-up question is whether this consistent performance still holds at a finer resolution, such as the angular resolution of the electromagnetic calorimeter (ECAL), $\Delta R_{E}\simeq 0.025$. This section addresses the expected AM performance at sub-HCAL angular resolution.
Before delving into the analysis of sub-HCAL angular resolution, we must pay attention to new features appearing at the smaller scales. The first important difference is that ECAL towers and tracks are implicitly distinguishable from HCAL towers in the EFlow object analysis at sub-HCAL angular resolution. If there is a constituent cluster with a size below $\Delta R_{H}$, that cluster is highly likely to consist of ECAL towers and tracks. Similarly, if a cluster is smaller than $\Delta R_{E}$, the cluster mostly consists of tracks.
This implicit accessibility of constituent type leads to an interesting consequence: particle identification (PID) information is also partially and implicitly accessible in sub-HCAL angular resolution analysis. Comparing energy deposits in the HCAL, ECAL, and tracks partially reveals how much energy is carried by certain types of particles. This is useful information when comparing jets with different PID compositions, and the use of the energy comparison has been studied in the context of jet flavor identification Park:2017rfb ; Nakai:2020kuu . Also, explicitly considering PID information improves jet tagging performances, and it has been demonstrated using particle level analysis, i.e., when there is no implicit utilization of angular resolution Qu:2019gqs .
Designing an analysis module and high-level features for capturing this constituent type information may further supercharge the tagging performance of AM. However, constructing a new module analyzing implicit PID information is beyond the scope of the paper. We will focus on tweaking the current AM setup to a finer resolution analysis and compare its performance to that of ParT.
For adjusting AM for higher resolution, we first have to be cautious about an inflated number of inputs. For example, the set of two-point energy correlations with trimming, $x_{\mathrm{trim}}$, described in Section 2.3, originally consisted of 60 inputs, i.e., (3 types of $S_{2,ab}$) $\times$ (20 bins). But at the ECAL angular resolution level, now we have to consider 80 bins per one $S_{2,ab}$, and it inflates the number of inputs by a factor of 4. Similarly, the set of MFs, $x_{\mathrm{MF}}$, described in Section 2.4, consists of 150 inputs, i.e., (3 Minkowski functionals) $\times$ (10 dilation scales) $\times$ (5 $p_{T}$ thresholds). Again, we have to consider four times more dilation scales to cover the same dilation scale range, so the total number of inputs will be inflated to 600.
The reason why we have to be cautious of inflated inputs is that it may make the training dataset more sparsely distributed on the input space. Jets in our dataset contain about 50 constituents, and the dimensionality of the input space of the AM for ECAL angular resolution analysis exceeds the dimension of the 4-momenta of jet constituents, $4\times 50$. In this case, less carefully tuned AM may require more training samples to compensate for the sparsity of the data.
Here, we present the results of AM tweaked for the ECAL angular resolution $\Delta R_{E}$. In this quick analysis, instead of using finer pixels and bins everywhere, we simply augment the previous AM for HCAL angular resolution analysis in order to avoid input dimension inflation, making the dataset sparse. Here is the list of changes:
  • Pixelated jet constituent multiplicity variables in $x_{\rm count}$, $x_{\rm subj}$ are now evaluated using jet images with a pixel size $\Delta R_{E}\times\Delta R_{E}$.
  • Two-point energy correlation analysis modules $\Phi^{\mathrm{trim}}$ and $\Phi^{\mathrm{trim}}$ additionally takes extra binned $S_{2,ab}$’s on a domain $[0,0.25)$ and with bin size 0.025.
  • Minkowski functional analysis module $\Phi^{\mathrm{MF}}$ additionally accounts Minkowski functionals with dilation scales in $[0.0125,0.125]$ and with step size 0.0125.
Namely, we count the sub-HCAL angular scale information up to the angular scale $R=0.25$, and we consider information above the ECAL scale only using inputs with HCAL angular resolution. With this setup, each of the analysis modules is able to take care of differences in their high-level inputs in ECAL and HCAL angular resolution around and below HCAL angular resolution. Also, 10 new inputs per each of $S_{2,ab}$ and MFs are considered so that the input dimensions are not inflated too much.
The classification performances of AM and ParT at this ECAL angular scale are shown in Table 7 and Table 8. Here, the average and uncertainties of the performance metrics are estimated using a bootstrap method. The ParT model used here is the same ParT model described in Section 3.2. The difference in the classification metrics between AM and ParT lies between $2\sigma$ to $5\sigma$ when we only count variation of classifier under change of random seeds. Again, the error in this comparison is mostly dominated by uncertainties of ParT training, and $\mathrm{P_{T}}$ vs. $\mathrm{H_{T}}$ shows the largest difference in metric about $5\sigma$. We will leave accounting for the additional high-level features for the smaller scales, further fine-tuning AM setup and binning of the inputs, and estimating the full uncertainty precisely for future studies.
AUC AUC AUC AUC
Full AM 0.9467$\pm$0.0002 0.9313$\pm$0.0001 0.608$\pm$0.004 0.618$\pm$0.001 0
ParT 0.9498$\pm$0.0011 0.9348$\pm$0.0007 0.620$\pm$0.005 0.649$\pm$0.006 0
Table 7: The average and standard deviations of AUC, considering ten networks trained using different random number seeds and trained on bootstrapped samples. Here, we use the same ParT model described in Section 3.2.
Model $\mathrm{P_{T}}$ vs. $\mathrm{P_{Q}}$ $\mathrm{H_{T}}$ vs. $\mathrm{H_{Q}}$
$R_{50\%}$ $R_{30\%}$ $R_{50\%}$ $R_{30\%}$
Full AM 096$\pm$1 392$\pm$04 67.5$\pm$0.4 281$\pm$06
ParT 106$\pm$4 430$\pm$26 73.9$\pm$2.0 302$\pm$13
Model $\mathrm{P_{Q}}$ vs. $\mathrm{H_{Q}}$ $\mathrm{P_{T}}$ vs. $\mathrm{H_{T}}$
$R_{50\%}$ $R_{30\%}$ $R_{50\%}$ $R_{30\%}$
Full AM 2.87$\pm$0.01 5.61$\pm$0.03 2.97$\pm$0.02 5.74$\pm$0.03
ParT 3.01$\pm$0.06 5.93$\pm$0.15 3.38$\pm$0.08 6.77$\pm$0.20
Table 8: The average and standard deviations of the rejection rates $R_{50\%}$ and $R_{30\%}$, considering ten networks trained using different random number seeds and trained on bootstrapped samples. Here, we use the same ParT model described in Section 3.2.